Jesteś tutaj

Observation

Status: 
Wykonane
Czas obserwacji: 
piątek, Kwiecień 13, 2012 - 20:30 to sobota, Kwiecień 14, 2012 - 19:00
Obserwator: 
Instruktor: 
Teleskop zaparkowany?: 
Tak
Kopuła zaparkowana?: 
Tak

Goto-targeting was off but now fixed.

A lot of humidity.

 

The photometric filter wheel is placed inside the CCD.

The spectrograph was left mounted on the telescope.